OpenStudy (anonymous):

find the inner product of a and b if a=<4,-2,-2> and b=<-7,-2,4>, and state whether the vectors are perpendicular

OpenStudy (anonymous):

so its -32?

OpenStudy (perl):

$$ \Large \bf \vec{a} ~\cdot \vec{b} = a1*b1 + a2*b2 + a3*b3 $$

OpenStudy (perl):

correct

OpenStudy (anonymous):

thanks and i have one more question

OpenStudy (anonymous):

and are they perpendicular?

OpenStudy (perl):

same deal :)

OpenStudy (anonymous):

so yes to perpendicular ?

OpenStudy (perl):

Two vectors are perpendicular iff the dot product is equal to zero. since it is not zero, not perpendicular

OpenStudy (anonymous):

ow thanks

OpenStudy (perl):

$$ \LARGE \bf \vec{a} ~\perp \vec{b} ~\Leftrightarrow~ \bf \vec{a} ~\cdot \vec{b} = 0 $$